OpenStudy (anonymous):

find exact values of the expression 2sin(pi/12)cos(pi/12)

OpenStudy (anonymous):

\[2\sin \frac{ \pi }{12 }\cos \frac{ \pi }{ 12 }\]

OpenStudy (sirm3d):

\[\sin \frac{ \pi }{ 6 }\] or 1/2 or 0.5

OpenStudy (anonymous):

hows you get that?

OpenStudy (sirm3d):

identity

OpenStudy (sirm3d):

\[\sin 2A=2\sin A \cos A\] \[A=\frac{ \pi }{ 12 }\] in your problem

OpenStudy (anonymous):

do i divide by 2?

OpenStudy (sirm3d):

the right side of the identity is your given, the left side will be the answer

OpenStudy (sirm3d):

use A=pi/12

OpenStudy (anonymous):

so its \[\frac{ 2 }{1 } \times \frac{ \pi }{ 12 }\]

OpenStudy (sirm3d):

that's pi/6 when simplified

OpenStudy (anonymous):

okay i got it, how would you simiplify tan pi/8

OpenStudy (sirm3d):

the property changes \[8\log_{8}30 \] to \[8\log_8 2 + 8\log_8 3 + 8\log_8 5\]

OpenStudy (sirm3d):

oops, wrong tab. hahaha

OpenStudy (anonymous):

ohh lol i was like what

OpenStudy (sirm3d):

\[\tan \left( \frac{ \pi }{ 8 } \right)=\tan \left( \frac{ 1 }{ 2 } *\frac{ \pi }{ 4 }\right)\] then use the identity for half-angle of tan

OpenStudy (anonymous):

the \[\tan \frac{ u }{ 2}= \frac{ 1-cosu }{ sinu } \]

OpenStudy (sirm3d):

no square on tangent?

OpenStudy (anonymous):

it doesnt have one in the book

OpenStudy (anonymous):

or do i use tan2x=2tanx/1-tan^2x

OpenStudy (sirm3d):

the first formula is correct. and use it

OpenStudy (anonymous):

how do distinguish between which one is sinu or cosu

OpenStudy (sirm3d):

\[\tan \frac{ \pi }{ 8 }=\tan \left( \frac{ \pi/4 }{ 2 } \right)\] use \[u=\frac{ \pi }{ 4 }\] on both sine and cos

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

after i get\[1-\frac{ \sqrt{2} }{ 2 }/\frac{ \sqrt{2} }{ 2 }\] what should i do?

OpenStudy (sirm3d):

rationalize the denominator by multiplying both sides by sqrt(2)

OpenStudy (sirm3d):

got to go. i see raised eyebrows. hahaha